Hello frnds..........i m yogita....preparing for iit jee-2012


 


Can anybody give me the proof for dearrangement series in permutations..........??????

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear yogita,

Any change in the existing order of things is called a derangement.

If 'n' things are arranged in a row, the number of ways in which they can, be deranged so that none of them occupies its original place is

 derangements                         

 and it is denoted by D(n).

A question on derangement can be of the following kind:

Illustration:

Supposing 4 letters are placed in 4 different envelopes. In how many ways can be they be taken out from their original envelopes and distributed among the 4 different envelopes so that no letter remains in its original envelope?

Solution:

Using the formula for the number of derangements that are possible out of 4 letters in 4 envelopes, we get the number of ways as :

 4!(1 - 1 + 1/2! - 1/3! + 1/4!) = 24(1 - 1 + 1/2 - 1/6 + 1/24) = 9.

 

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi